You come across a small office building for sale with an asking price of $400,000. After careful analysis, you estimate that you’ll receive $46,000 in annual profit before financing.

For each part of this question, show your work

a). What is your projected Return-on-Assets for this investment?  

b). After talking with your lender, you find you can easily borrow 60% of the property’s purchase price. Interest and fees related to the loan will cost you $14,000 per year.

What is your projected Return-on-Equity for this investment?

Show your Annual Profit After Financing calculation by constructing a table

c). Being a savvy investor, you want to find out what would happen if there is downturn in the market. In the event of a market downturn, you estimate the annual profit before financing would be negative $10,000.

What would the Return-on-Assets for this investment be in the event the market downturn above?

d). Under the same assumptions as Q3(c), what would the Return-on-Equity for this investment be?

Show your Annual Profit After Financing calculation by constructing a table as seen in the “Risks of Financial Leverage” discussion in lecture.

e). What do the two cases above tell us about the effect of leverage on real estate investments?
Answer #1

A)

Item Values Answer
Estimated Income 46,000
Total Investments 4,00,000
% Projected Return on Investments 46,000/4,00,000 11.50%

B)

Item Values Answer
Estimated Income 46,000
Interest Expenses 14,000
Net Income post interest Expenses (Annual Profit) 46,000-14,000 = 32,000
Equity 40% of total investments 4,00,000*40%= 1,60,000
Total Investments 4,00,000
% Return on Equity 32,000/1,60,000 20%

C)

Item Values Answer
Annual Income Before financing -10000
Interest Expenses 14000
Net Income post interest Expenses (Annual Profit) -10000-14000 = -24000
Total Investments 400,000.00
% Return on Assets -24,000/4,00,000 6%

D)

Items Values Answer
Total Investments 400,000.00
Debt 60% 240,000.00
Equity 40% 160,000.00
Annual Income Before financing -10,000.00
Interest Expenses 14,000.00
Net Income post interest Expenses (Annual Profit) -10,000-14,000 = -24,000.00
Return on Equity -24,000/160000 -15%

E) Financial leverage is 2 edged sword. Financial leverage both increases the potential returns to common stockholders and the risk. In absence of financial leverage Return on Assets(ROA) will be equal to return on Equity(ROE). As seen in a & b part due to financial leverage ROE has increased from 11.5% to 20%. Also when profit turns into loss ROE turns into deep -ve from -6.00%(in absence of financial leverage) to -15%. So financial leverage increase both risk and return.